OpenStudy (anonymous):

Need help solving this equation... (k * Q1 * Q)/(x^2) = (k * 4Q1 * Q)/(L-x^2)

OpenStudy (anonymous):

Q1 and Q are constant, so is k. Variables are L and x.

Parth (parthkohli):

Right, you should cancel them off from both sides.

OpenStudy (anonymous):

so right off the bat I can cancel all of those and will be left with.... 1/x^2 = 4 / (L-x)^2 This was supposed to be (L-x)^2 in problem description.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Hi :) Cross multiply!

OpenStudy (anonymous):

Hi, so then I will have... 4x^2 = (L-x)^2

OpenStudy (anonymous):

Correct! There's no way to solve this equation until you're specified anything else.

OpenStudy (anonymous):

then I can expand (L-x)^2 to L^2 -2Lx + x^2

OpenStudy (anonymous):

Yes, very much!

OpenStudy (anonymous):

how about L = x + ( L -x) ?

OpenStudy (anonymous):

Right! But can you explain how'll use this further?

OpenStudy (anonymous):

No. My book has a solution for this....I have been messing with this for awhile and agree with you that there is no way to solve it wihtout something I am missing. Ill carefully look at the problem again.

OpenStudy (anonymous):

Can you write what your book says here?

OpenStudy (anonymous):

Two charges -Q1 and -4Q1 are a distance L apart. These 2 charges are free to move but do not because a third charge holds the two in equilibrium. What must be the magnitude and placement of the third charge?

OpenStudy (anonymous):

Hey wait! I think the question was changed

OpenStudy (anonymous):

Oh alright! Principle of superposition!

OpenStudy (anonymous):

Yes Thats where the equations I was working with came from.

OpenStudy (anonymous):

Thats not (L - x)^2 but L-x^2 right?

OpenStudy (anonymous):

In that case I think its better

OpenStudy (anonymous):

It should be (L-x)^2

OpenStudy (anonymous):

Oh ok. Ive got a little rusty at my electrostatics.

OpenStudy (anonymous):

I think you should ask this in physics! Those folks are good at this

OpenStudy (anonymous):

Its just a math problem at this point.

OpenStudy (anonymous):

I understand what you mean

OpenStudy (anonymous):

Aravind any inputs?

OpenStudy (aravindg):

One thing we can directly say is that the third charge is positive. (k * Q1 * Q)/(x^2) = (k * 4Q1 * Q)/(L-x)^2

OpenStudy (aravindg):

You can cancel k*Q*Q1 on both sides. Then take square root :)

OpenStudy (anonymous):

How could it be so simple lol.............so could you have solved it if you cross multiplied and kept going that route?

OpenStudy (aravindg):

Why go the complicated way when you can do it in the simple way? Yes ultimately in any method we get the same answer.

OpenStudy (anonymous):

Oh yes I forgot about the square root! That was very stupid lol

OpenStudy (aravindg):

Yes all the terms come in a way that compels us to take the root. 4,x^2,(l-x)^2

OpenStudy (anonymous):

Thank you!